Joplin Corporation produces syrups that it sells to candy makers. On November 1, it had no work-in-process inventory. It started production of 46,200 gallons of syrup in November and completed production of 44,000 gallons. The costs of the resources used by Joplin in November consist of the following:

Materials $ 64,604
Conversion costs (labor and overhead) 53,800


Required:

The production supervisor estimates that the ending work-in-process is 70 percent complete on November 30. Compute the cost of syrup completed and the cost of the syrup in work-in-process ending inventory as of November 30. (Do not round intermediate calculations.)
Answer #1

Ending work in process = 46200-44000 = 2200 gallon

Equivalent unit = 44000+(2200*70%) = 45540 gallon

Cost per equivalent unit = (64604+53800)/45540 = 2.6

Cost of syrup completed = 2.6*44000 = $114400

Cost of syrup in ending inventory work in process = 1540*2.6 = $4004